# Quantum accessible information and classical entropy inequalities

### v4

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. We show that the recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
@@ 2 Accessible information: a criterion for optimality
$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$ $\longrightarrow$ $\min$ (11)
@@
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$ (15)
@@
$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (16)
@@
$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (17)
@@ 3 Example: two pure states
$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$ (22)
@@
$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$ (23)
@@
$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$ (25)
@@
Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (8) which via (24), (26) amounts to the following parametric family of tight lower bounds for the classical binary entropy  [^3] See the dynamical plot of the inequality at https://www.desmos.com/Calculator/11ikk1i8fb
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$ (35)
@@ 4.2 The conjecture: “quantum pyramids”
The cosine between the edges is $\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}\,,\,j\neq k,$ thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute, $r_{0}=r_{1}\,(b=0)$ – to the orthogonal, and $r_{0}<r_{1}\,(b<0)$ – to the obtuse pyramids. The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle,$ and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat pyramid.
@@
$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$ (51)
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (53)
@@
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (63)
@@ 4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
@@
The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (37) with $r_{0}<r_{1}$ $(b<0)$ . The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
The suggestion for a proof of the inequalities is given in section 5.2.
@@ 4.5 The case of flat pyramid
The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the average state of the ensemble $\mathcal{E}$
$\overline{\rho}=\frac{1}{m-1}\left(I-|0\rangle\langle 0|\right)$
@@
where $|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle.$ Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
@@
Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$, hence the second equality follows.
@@
For $m=3$ the inequality (83) reduces to (66). Indeed, as shown below, in (83) one can restrict to real $z_{j}$. Then $|\psi\rangle$ is a vector in the real plane $\langle e_{0}|\psi\rangle=0$ and
$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\|\tilde{\psi}_{j}\|\,\cos\left(\alpha+\frac{2\pi j}{3}\right),\quad j=1,2,3,$
where $\||\tilde{\psi}_{j}\rangle\|=\||e_{j}\rangle-\sqrt{1/3}|e_{0}\rangle\|=\sqrt{2/3}$, hence (83) is the same as (66).
@@ 5.1 The case of theorem 2
$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$ (87)
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (88)
@@ 5.2 The case of theorem 3
$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl|\sum\limits_{j=1}^{m}z_{j}\Bigl|^{2}+\tilde{\mu}_{1}(p)\geq 0.$ (92)
@@
Suggestion for proof.
As it has been established in lemma 1, a real case is enough for analysis.
For real vectors $(z_{j})_{j=1}^{m}$ from the unit sphere let us consider the Lagrange function
@@
The necessary condition is
@@
and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since the equation
@@
for a variable $z$ has at most three real solutions. The fact the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then three real roots was actually used in proof of lemma 2 for equation (86) which is similar to (93) with $\leq 2$ roots. However, here we consider the entire real line, so the maximum number of real roots is three.
Consider the following cases.
@@
Then $\lambda+\log e=0$ and the equation (93) has roots 0 and $\pm s$. Therefore the minimizer of the function $\tilde{F}(z)$ should have the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$. The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl[\dfrac{m-1}{m},1\Bigr)$. The derivative is
@@
Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
@@
because of (79). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
@@
To reduce the number of different values to two, we will introduce a new Lagrange function $\hat{L}$. The critical points of $L$ are related with critical points of the Lagrange function
$\hat{L}(s_{1},s_{2},s_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}\log s_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}k_{j}s_{j}\Bigl)^{2}$-
$-\lambda\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}-1\Bigl)-\nu\Bigl(\sum\limits_{j=1}^{3}k_{j}-m\Bigl).$
New necessary conditions are
$-\dfrac{1}{2}\dfrac{\partial\hat{L}}{\partial s_{l}}=k_{l}s_{l}\log s_{l}^{2}+\tilde{\mu}_{0}(p)k_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+(\lambda+\log e)k_{l}s_{l}=0,$
$-\dfrac{\partial\hat{L}}{\partial k_{l}}=s_{l}^{2}\log s_{l}^{2}+2\tilde{\mu}_{0}(p)s_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\lambda s_{l}^{2}+\nu=0$
for all $1\leq l\leq 3$.
Thus, we have
$\dfrac{1}{2}s_{l}\dfrac{\partial\hat{L}}{\partial s_{l}}+k_{l}\dfrac{\partial\hat{L}}{\partial k_{l}}=k_{l}s_{l}\Bigl(s_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)-\nu k_{l}=0.$
Case 2.1. $k_{1},k_{2},k_{3}\neq 0$.
For all $1\leq l\leq 3$ it holds $s_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\nu$, which can take place, if the set $\{s_{1},s_{2},s_{3}\}$ has no more than two different values (which is realized, formally, for fractional multiplicities). This can be regarded as the next case.
Case 2.2. $k_{3}=0$.
Up to permutation and the global sign, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}>0$}, $s_{2}<0$}. The components of the minimizer should have different signs; otherwise, changing the sign of a single component would yield a lower value for $\tilde{F}$.
@@
$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl(\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl)^{2}-\tilde{\mu}_{1}(p,1/m),$ (94)
and the statement follows from lemma 5.

### v5

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. We show that the recently obtained optimality criterion (A.S. Holevo, Lobachevskii J. Math., 43:7 (2022), 1646-1650), when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) (B.-G. Englert and J. Řeháček, J. Mod. Optics 57 N3 (2010) 218-226) is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also an approach to the proof of the conjectures concerning globally information-optimal observables for quantum pyramids.
Keywords and phrases: quantum state ensemble, accessible information, convex optimization, quantum pyramid, tight lower bound for Shannon entropy.
@@ 2 Accessible information: a criterion for optimality
$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$ $\longrightarrow$ $\min$
@@
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$
@@
$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$
@@
$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$
@@ 3 Example: two pure states
$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$
@@
$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$
@@
$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$
@@
Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (8) which via (17), (18) amounts to the following parametric family of tight lower bounds for the classical binary entropy
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$
@@ 4.2 The conjecture: “quantum pyramids”
The cosine of angle between any two different edges is
$\xi\equiv\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}=2ab+(m-2)b^{2}\,,\quad\,j\neq k,$ (33)
thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute ($0<\xi<1$), $r_{0}=r_{1}\,(b=0)$ – to the orthogonal ($\xi=0$), and $r_{0}<r_{1}\,(b<0)$ – to the obtuse $\left(-\frac{1}{\left(m-1\right)}<\xi<0\right)$ pyramids. The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle$ ($\xi=1$), and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat $\left(\xi=-\frac{1}{\left(m-1\right)}\right)$ pyramid.
@@
$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$
@@
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$
@@ 4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first briefly survey the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ In the section 4.5 we perform the analysis based on our optimality criterion showing that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
@@
A generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (28) with $r_{0}<r_{1}$ $(b<0)$ . The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
The suggestion for a proof of the inequalities combining analytical argument with computer-aided reasoning is given in section 5.2.
@@ 4.5 The case of flat pyramid
The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the edges (28) of the pyramid
$|\psi_{j}\rangle=\sqrt{\frac{m-1}{m}}|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle;\quad j=1,\dots,m$
all lie in the hyperplane $\mathcal{H}_{0}=\left\{\psi:\langle e_{0}|\psi\rangle=0\right\}$. The vectors are linearly dependent and their convex hull is a regular symplex in $\mathcal{H}_{0}.$ The average state of the ensemble $\mathcal{E}$
$\overline{\rho}=\frac{1}{m-1}\left(I-|e_{0}\rangle\langle e_{0}|\right)$
@@
where
$|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle=\sqrt{\frac{m-1}{m}}|\psi_{j}\rangle.$ (69)
Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
@@
Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$i equality follows.
@@
For $m=3$ the inequality (71) reduces to (53). Indeed, as shown in lemma 1, one can restrict to real $z_{j}$ in (71). Then $|\psi\rangle$ is a vector in the real plane $\mathcal{H}_{0}$ and by (69)
$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\sqrt{2/3}\langle\psi_{j}|\psi\rangle=\sqrt{2/3}\cos\alpha_{j},\quad j=1,2,3,$
where $\alpha_{j}$ is the angle between the vector $|\psi\rangle$ and the ensemble vector $|\psi_{j}\rangle$ in the plane $\mathcal{H}_{0}.$ The angle between any two different vectors of the ensemble is $2\pi/3$ because $\langle\psi_{j}|\psi_{k}\rangle=-1/2$ according to (33). Therefore $\alpha_{j}=\alpha+\frac{2\pi j}{3},\quad j=1,2,3,$ thus we come to (53).
Let us also explain how (53) follows from the result of the paper [14]. A simple calculation shows that (53) is equivalent to
$\log 3-\frac{1}{3}\sum_{j=1}^{3}\left(1+\cos\left(2\alpha+\frac{4\pi j}{3}\right)\right)\log\left(1+\cos\left(2\alpha+\frac{4\pi j}{3}\right)\right)\geq\log 2;$
A special case of lemma 3 in [14] (with $M=3$) implies that the deductible in the lefthand side is maximized for $\alpha=\pi/2+\pi k$ (giving in fact the accessible information $\log\left(3/2\right)$). The proof in [14] is based on the decomposition of the function $\left(1+x\right)\log\left(1+x\right)$ into the power series and a clever manipulation with the terms of the decomposition.
@@ 5.1 The case of theorem 2
$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$
@@ 5.2 The case of theorem 3
$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl|\sum\limits_{j=1}^{m}z_{j}\Bigl|^{2}+\tilde{\mu}_{1}(p)\geq 0.$
@@
Suggestion for proof with computer verification.
As it has been established in lemma 1, it suffices to consider real vectors.
For real vectors $(z_{j})_{j=1}^{m}$ on the unit sphere let us consider the Lagrange function
@@
The first-order conditions yield
@@
and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since equation
@@
in a variable $z$ has at most three real solutions. The fact that the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more than three real roots was actually used in the proof of lemma 2 for equation (74) which is similar to (78) with $\leq 2$ roots. However, here we consider the entire real line, so the maximum number of real roots is three.
We consider two cases.
@@
Then $\lambda+\log e=0$ and the equation (78) has roots 0 and $\pm s$. Therefore the minimizer of $\tilde{F}(z)$ must be of the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$. The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
First, we prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl[\dfrac{m-1}{m},1\Bigr)$. The derivative is
@@
Here, $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, hence $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
@@
by (66). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
@@
The Lagrange multipliers method predicts that the minimizer has at most 3 different values, that are in our case nonzero. They are denoted as $s_{1}<0<s_{2}\leq s_{3}$ (without loss of generality we can assume that only $s_{1}$ is negative) with multipliers $k_{1},k_{2},k_{3}$, such that $k_{1}+k_{2}+k_{3}=m$ and $k_{1}s_{1}^{2}+k_{2}s_{2}^{2}+k_{3}s_{3}^{2}=1$. Our hypothesis is that, however, there is only two distinct nonzero values among $s_{1},s_{2},s_{3}$. This fact can be verified by computer experiments; we performed numerical simulations for $m$ from $3$ to $50$ and for the parameter $a=\tilde{\mu}_{0}(p)>0$ taking values in $\{0.1,0.5,1,5,10,100,\infty\}$ (see Appendix).
Thus, up to permutation and an overall sign change, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}<0$}, $s_{2}>0$}. The components of the minimizer must have opposite signs; otherwise, changing the sign of a single component would decrease $\tilde{F}$.
@@
$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl(\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl)^{2}-\tilde{\mu}_{1}(p,1/m),$
and the result follows from lemma 5.